OpenStudy (anonymous):

What is the slope of the line passing through (1, 2) and (3, 8)? slope = 1/7 slope = 1/3 slope = 3 slope = 7 i got 1/3 wbu

OpenStudy (anonymous):

slope formula is \[\frac{ y2-y1 }{ x2-x1 }\] \[\frac{ 8-2 }{ 3-1 }\]

OpenStudy (anonymous):

i got 6 1/3 how is tht

OpenStudy (anonymous):

I get 6/2=3/1=3

OpenStudy (anonymous):

8-2=6 3-1=2 6/2=3/1=3

OpenStudy (anonymous):

so whats the awnser heather

OpenStudy (anonymous):

3

OpenStudy (anonymous):

tyty

OpenStudy (anonymous):

yup np :)